Find all real or imiganary soultions to the equation:
√7 + 29 = x + 3
Thanks for your help.

Answer by RAY100(1637)

sqrt7 +29 = x +3
.
sqrt 7 +29 -3 =x
.
+/-(2.646) +29-3 = x
.
23.354,, 28.646 =x
.
check (23.354) ,,,, -sqrt7+29= 23.354 +3 = 26.354,,,,ok
.
(28.646) ,,,,+sqrt7 +29 = 28.646 +3 =31.646,,,,,,ok
.